# 3S-TSE: Efficient Three-Stage Target Speaker Extraction for Real-Time and Low-Resource Applications

### 18 Dec 2023 v1

@@ 1 Introduction
Unlike traditional speech enhancement systems, target speaker extraction emerges as a highly practical solution to the well-known cocktail party problem [1, 2]. In this approach, the system is designed to exclusively extract the speech of a designated target speaker from a mixture of voices, using the reference speech of that target speaker as a guide. This task holds a pivotal role in various applications, including voice recognition, multi-speaker transcription, and assistive hearing technologies [3]. It effectively amplifies the voice of the target speaker while concurrently suppressing interfering voices, thereby enhancing the clarity and intelligibility of the desired speech signal. Inspired by human auditory attention, current deep learning speaker extraction methods [4, 5, 6, 7, 8] predominantly employ a two-subnet architecture comprising a speaker encoder and speech extractor [9, 10, 11, 12, 13, 14]. The speaker encoder models the voiceprint of the target speaker and subsequently guides the speech extractor to isolate the speech signal associated with that specific speaker. However, the voiceprint extracted by the encoder and the mixture signal features occupy disparate high-dimensional spaces. Bridging this gap requires extensive end-to-end neural networks, which are often complex and large [15, 12].
@@
Unfortunately, the trend in literature is to develop increasingly large and complex models in pursuit of higher accuracy and precision, often at the expense of computational efficiency. This focus on performance has largely overshadowed the need for models that can operate effectively under real-time, low-resource conditions, such as those required by edge computing devices. To the best of our knowledge, there exists no prior work addressing the problem of target speaker extraction under real-time and resource-constrained conditions, such as those encountered in edge computing environments.
@@ 2.4 Second Stage: GSC
Griffiths introduced the GSC, which adapts the linearly constrained minimum-variance proposed by Frost [20] to an unconstrained scenario. $[y_{1}(n),y_{2}(n),\ldots,y_{M}(n)]$ represents the speech inputs from the M microphones. The GSC structure primarily comprises two subsystems: 1. The upper branch houses the fixed beamformer, characterized by a delay alignment structure and a fixed weight set $W_{c}=\left[w_{c1},w_{c2},\ldots,w_{cM}\right]$. By manipulating $W_{c}$, one can control both the gain and beam width of this beamformer. 2. The lower branch contains the blocking matrix $B$, responsible for noise component estimation. The outputs from these two subsystems undergo adaptive cancellation via an adaptive canceller paired with a coefficient matrix $W_{a}$, producing the enhanced speech. Griffiths used the traditional least mean square (LMS) algorithm to derive the adaptive weights $W_{a}$. Here, $y_{a}(n)$ is the fixed beamformer output, $y_{z}(n)$ denotes the noise in $y_{a}(n)$, and $y_{o}(n)$ represents the enhanced speech.
The output angle from the TS-DOA module is utilized as the target angle in GSC for computing the fixed beamforming weight $W_{a}$. The output of GSC is sent to the subsequent Post-Processing module as the preliminary target speaker extraction result.
@@ 3.1 Dataset and setup
As we are performing online per-frame DOA estimation of the target speaker, silence segments are labeled without a direction. This results in 37 possible output targets per frame: 36 direction classes and a 1 no-direction silence class. Voice activity detection (VAD) [25] is applied on the clean target speech to differentiate silent frames from voiced speech.
@@ 4 CONCLUSIONS
Acknowledgments: This research was supported by the China National Nature Science Foundation (No. 61876214).

### 05 Jan 2024 v2

@@ 1 Introduction
Unlike traditional speech enhancement systems, target speaker extraction emerges as a highly practical solution to the well-known cocktail party problem [1]. In this approach, the system is designed to exclusively extract the speech of a designated target speaker from a mixture of voices, using the reference speech of that target speaker as a guide. This task holds a pivotal role in various applications, including voice recognition, multi-speaker transcription, and assistive hearing technologies [2]. It effectively amplifies the voice of the target speaker while concurrently suppressing interfering voices, thereby enhancing the clarity and intelligibility of the desired speech signal. Inspired by human auditory attention, current deep learning speaker extraction methods [3, 4, 5, 6, 7] predominantly employ a two-subnet architecture comprising a speaker encoder and speech extractor [8, 9, 10, 11, 12]. The speaker encoder models the voiceprint of the target speaker and subsequently guides the speech extractor to isolate the speech signal associated with that specific speaker. However, the voiceprint extracted by the encoder and the mixture signal features occupy disparate high-dimensional spaces. Bridging this gap requires extensive end-to-end neural networks, which are often complex and large [13, 10].
@@
Unfortunately, the trend in literature is to develop increasingly large and complex models in pursuit of higher accuracy and precision, often at the expense of computational efficiency. This focus on performance has overshadowed the need for models that can operate effectively under real-time, low-resource conditions [18], such as those required by edge computing devices. To the best of our knowledge, there exists no prior work addressing the problem of target speaker extraction under real-time and resource-constrained conditions, such as those encountered in edge computing environments.
@@ 2.4 Second Stage: GSC
Griffiths introduced the GSC, which adapts the linearly constrained minimum-variance proposed by Frost [19] to an unconstrained scenario. $[y_{1}(n),y_{2}(n),\ldots,y_{M}(n)]$ represents the speech inputs from the M microphones. The GSC [21] structure primarily comprises two subsystems: 1. The upper branch houses the beamformer, characterized by a delay alignment structure and a weight set $W_{c}=\left[w_{c1},w_{c2},\ldots,w_{cM}\right]$. By manipulating $W_{c}$, one can control both the gain and beam width of this beamformer. 2. The lower branch contains the blocking matrix $B$, responsible for noise component estimation. The outputs from these two subsystems undergo adaptive cancellation via an adaptive canceller paired with a coefficient matrix $W_{a}$, producing the enhanced speech. Griffiths used the traditional least mean square (LMS) algorithm to derive the adaptive weights $W_{a}$. Here, $y_{a}(n)$ is the beamformer output, $y_{z}(n)$ denotes the noise in $y_{a}(n)$, and $y_{o}(n)$ represents the enhanced speech.
The output angle from the TS-DOA module is utilized as the target angle in GSC for computing the beamforming weight $W_{a}$. The output of GSC is sent to the subsequent Post-Processing module as the preliminary target speaker extraction result.
@@ 3.1 Dataset and setup
As we are performing online per-frame DOA estimation of the target speaker, silence segments are labeled without a direction. This results in 37 possible output targets per frame: 36 direction classes and a 1 no-direction silence class. Voice activity detection (VAD) is applied on the clean target speech to differentiate silent frames from voiced speech.
@@ 4 CONCLUSIONS
Acknowledgments: This research was supported by the China National Nature Science Foundation (No. 61876214). This work was also supported by the Open Fund (KF-2022-07-009) of Key Laboratory of Urban Land Resources Monitoring and Simulation, Ministry of Natural Resources, China.